OpenStudy (anonymous):

Translate into an equation and solve. The sum of a number, n, and seven is twenty less than four times the number.

OpenStudy (anonymous):

3n-27=0 n=27/3

myininaya (myininaya):

sum n + 7 is = twenty less than four times n 4n-20

myininaya (myininaya):

\[n+7=4n-20\]

OpenStudy (anonymous):

what is N? I don't understand that part

myininaya (myininaya):

solve for n

myininaya (myininaya):

amu already did it

myininaya (myininaya):

you can simplify his fraction

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

n is a simple unknown valued number, which we have to find